OpenStudy (anonymous):

write a polynomial function in standard form with given zeros. x= 0, 0, 2

OpenStudy (triciaal):

one approach with given zeros. x= 0, 0, 2 factors (x-0)(x - 0)(x-2) = 0 x^2(x-2)= 0 x^3 -2 x^2 f(x) = x^3 -2x^2